a invest 10% more than b, b invest 10% less than c, what did c invest if total investement is 5780?

Answer by checkley79(3341)

A=1.1B
B=.9C
A+B=C=5780
1.1B+.9C+C=5780
1.1(.9C)+.9C+C=5780
.99C+.9C+C=5780
2.89C=5780
C=5780/2.89
C=$2000 AMOUNT C INVESTED.
PROOF:
B=.9*2000=1800 B INVESTMENT.
A=1.1*1800=1980 A INVESTMENT.
1980+1890+2000=5780
5780=5780